# VegasAfterglow: A High-Performance Framework for Gamma-Ray Burst Afterglows

### v1

@@ 6.4 Radiative Fireball
As discussed in Section 3.2.3, when the electron energy fraction $\epsilon_{e}$ is high (e.g., $\sim 1$) and the blast wave enters the fast-cooling regime ($\nu_{c}<\nu_{m}$), the system becomes radiative. In this case, the shock dynamics deviate significantly from the adiabatic scenario (where $\epsilon_{\rm rad}=0$), which is commonly assumed in most afterglow models. Fig. 13 presents the optical and X-ray light curves for a radiative blast wave. Compared to the adiabatic case, the emission is markedly fainter and exhibits an earlier jet break, primarily due to the more rapid decrease in the Lorentz factor caused by radiative energy losses.
@@ 6.5 Reverse Shock Emission
Fig. 14 presents examples of on-axis optical ($10^{14}$ Hz) light curves from the forward and reverse shocks under different top-hat jet injection durations (with identical other parameters). The blue dashed lines represent the forward shock emission, while the red dashed lines correspond to the reverse shock. When the injection duration $\tau$ is much shorter than the deceleration time (e.g., $0.1$ s and $1$ s in this case), the reverse shock operates in the thin-shell regime. In this regime, the reverse shock crosses the entire ejecta shell around the deceleration time and reaches the same peak $\Gamma_{43}\sim 1.1$, indicated in the upper right panel of Figure 3. This results in nearly identical reverse shock light curves, as seen in the upper panels. The reverse shock emission scaling deviates from its $\propto t^{5}$ dependence near the shock crossing time as the system transitions out of the deep Newtonian regime. This transition causes the Lorentz factors $\Gamma_{3}$ and $\Gamma_{4}$ to diverge. As a result, the scaling of the relative Lorentz factor, $\Gamma_{43}$-1, also fails to follow the derived $\propto t^{3}$ relation, as demonstrated in the upper right panel of Figure 3.
@@ 6.6 Energy Injection
Fig. 15 shows an example of on-axis optical light curves of energy injection to a tophat jet with different total injection energies, mimicking the same injection function as magnetar spin-down as described in Equation 10. We fix the injection duration to be $10^{5}$ s for better visual comparison. A plateau feature is formed as the total injected energy is much larger than the initial kinetic energy of the jet.
@@ 6.7 User-defined jet
Fig. 16 illustrates a random example of a user-defined jet where the energy and Lorentz factor profiles are specified by Python functions. The upper panel presents an example using a damped cosine function for the energy profile and a damped sine function for the Lorentz factor profile. The bottom panel displays the corresponding optical ($10^{14}$ Hz) light curves calculated for several viewing angles. The VegasAfterglow code supports any arbitrary functions of $\phi$ and $\theta$ for energy, Lorentz factor, magnetization, and arbitrary functions of $\phi$, $\theta$, and $t$ for energy injection and matter injection, as detailed in Section 2.1.
@@ 7.2 Scientific Findings
    We derive, for the first time, a general analytical shock jump condition valid for arbitrary upstream magnetization $\sigma_{\rm u}$ and relative Lorentz factor $\Gamma_{\rm ud}$. The key result, providing the downstream four-velocity, is presented in Equation 26. In the limit $\sigma_{\rm u}\to 0$, this solution reduces to the classical Blandford-McKee result, while in the high-$\sigma$, highly relativistic regime, it asymptotically approaches the Kennel-Coroniti solution. This general formulation enables consistent modeling of both forward and reverse shocks across the full range of magnetization and Lorentz factor regimes.
@@
    Our simulations of detailed reverse shock crossing dynamics reveal that, in the thin-shell regime, reverse shocks are systematically weaker than previously estimated by analytical approximation in the literature. Specifically, we find that the peak relative Lorentz factor is typically $\Gamma_{43}\sim 1.1$, significantly lower than the commonly cited value of $7/4$. This suggests that predictions of reverse shock emission based on earlier analytical scalings, which rely on $\Gamma_{43}-1$, are likely overestimated.
@@ Acknowledgements
We thank all contributors who helped improve VegasAfterglow during its development. In particular, we are grateful to Weihua Lei, Shaoyu Fu, Iris Yin, Cuiyuan Dai, Liang-Jun Chen and Binbin Zhang for their invaluable efforts as beta testers, providing constructive feedback and assisting with bug fixes.

### v3

@@ 6.4 Radiative Fireball
As discussed in Section 3.2.3, when the electron energy fraction $\epsilon_{e}$ is high (e.g., $\sim 1$) and the blast wave enters the fast-cooling regime ($\nu_{c}<\nu_{m}$), the system becomes radiative. In this case, the shock dynamics deviate significantly from the adiabatic scenario (where $\epsilon_{\rm rad}=0$), which is commonly assumed in most afterglow models. Figure 13 presents the optical and X-ray light curves for a radiative blast wave. Compared to the adiabatic case, the emission is markedly fainter and exhibits an earlier jet break, primarily due to the more rapid decrease in the Lorentz factor caused by radiative energy losses.
@@ 6.5 Reverse Shock Emission
Figure 14 presents examples of on-axis optical ($10^{14}$ Hz) light curves from the forward and reverse shocks under different top-hat jet injection durations (with identical other parameters). The blue dashed lines represent the forward shock emission, while the red dashed lines correspond to the reverse shock. When the injection duration $\tau$ is much shorter than the deceleration time (e.g., $0.1$ s and $1$ s in this case), the reverse shock operates in the thin-shell regime. In this regime, the reverse shock crosses the entire ejecta shell around the deceleration time and reaches the same peak $\Gamma_{43}\sim 1.1$, indicated in the upper right panel of Figure 3. This results in nearly identical reverse shock light curves, as seen in the upper panels. The reverse shock emission scaling deviates from its $\propto t^{5}$ dependence near the shock crossing time as the system transitions out of the deep Newtonian regime. This transition causes the Lorentz factors $\Gamma_{3}$ and $\Gamma_{4}$ to diverge. As a result, the scaling of the relative Lorentz factor, $\Gamma_{43}$-1, also fails to follow the derived $\propto t^{3}$ relation, as demonstrated in the upper right panel of Figure 3.
@@ 6.6 Energy Injection
Figure 15 shows an example of on-axis optical light curves of energy injection to a tophat jet with different total injection energies, mimicking the same injection function as magnetar spin-down as described in Equation 10. We fix the injection duration to be $10^{5}$ s for better visual comparison. A plateau feature is formed as the total injected energy is much larger than the initial kinetic energy of the jet.
@@ 6.7 User-defined jet
Figure 16 illustrates a random example of a user-defined jet where the energy and Lorentz factor profiles are specified by Python functions. The upper panel presents an example using a damped cosine function for the energy profile and a damped sine function for the Lorentz factor profile. The bottom panel displays the corresponding optical ($10^{14}$ Hz) light curves calculated for several viewing angles. The VegasAfterglow code supports any arbitrary functions of $\phi$ and $\theta$ for energy, Lorentz factor, magnetization, and arbitrary functions of $\phi$, $\theta$, and $t$ for energy injection and matter injection, as detailed in Section 2.1.
@@ 7.2 Scientific Findings
    We derive, for the first time, a general analytical shock jump condition valid for arbitrary upstream magnetization $\sigma_{\rm u}$ and relative Lorentz factor $\Gamma_{\rm ud}$. The key result, providing the downstream four-velocity, is presented in Equation 26. In the limit $\sigma_{\rm u}\to 0$, this solution reduces to the classical Blandford-McKee result, while in the $\sigma, high$ly relativistic regime, it asymptotically approaches the Kennel-Coroniti solution. This general formulation enables consistent modeling of both forward and reverse shocks across the full range of magnetization and Lorentz factor regimes.
@@
    Our simulations of detailed reverse shock crossing dynamics reveal that, in the thin-shell regime, reverse shocks are systematically weaker than previously estimated by analytical approximation in the literature. Specifically, we find that the peak relative Lorentz factor is typically $\Gamma_{43}\sim 1.02$, significantly lower than the commonly cited value of $7/4$. This suggests that predictions of reverse shock emission based on earlier analytical scalings, which rely on $\Gamma_{43}-1$, are likely overestimated.
@@ Acknowledgements
We thank all contributors who helped improve VegasAfterglow during its development. In particular, we are grateful to Weihua Lei, Shaoyu Fu, Iris Yin, Cuiyuan Dai, Liang-Jun Chen and Binbin Zhang for their invaluable efforts as beta testers, providing constructive feedback and assisting with bug fixes. The authors acknowledge NASA 80NSSC23M0104, NASA 80NSSC20M0043, the Nevada Center for Astrophysics for support.